Could you please help me solve this system of linear equations?
It's -2x+y=1.3 and 2(0.5x-y)=4.6

Answer by Cromlix(4381)

Hi there,
-2x+y=1.3 and 2(0.5x-y)=4.6
I will use elimination method
for simultaneous equations.
- 2x + y = 1.3........(1)
2(0.5x - y) = 4.6.....(2)
Multiply out Equation (2)
- 2x + y = 1.3........(1)
x - 2y = 4.6..........(2)
Multiply Equation (2) by 2
- 2x + y = 1.3........(1)
2x - 4y = 9.2.........(2)
Add Equations (1) + (2)
....-3y = 10.5
.....y = -3.5
Substitute y = -3.5 into
Equation (1)
- 2x + (-3.5) = 1.3........(1)
-2x - 3.5 = 1.3............(10
-2x = 1.3 + 3.5
-2x = 4.8
x = -2.4
{x = -2.4 & y = -3.5}
Hope this helps :-)